# Transition to Petschek Reconnection in Subrelativistic Pair Plasmas: Implications for Particle Acceleration

### v1

@@ 1 Introduction
Reconnection is a ubiquitous plasma process that can rapidly convert magnetic energy into particle energy via rearrangement of the magnetic field topology. Originally posited to explain solar flares, reconnection now plays a key role in the theories of a multitude of astrophysical systems, from the heliosphere to compact objects. Observations and simulations suggest that reconnection can energize particles far beyond thermal energies and can generate hard power-law energy spectra (L. Sironi & A. Spitkovsky, 2014; F. Guo et al., 2014). The conditions under which this particle acceleration occurs, the precise mechanisms responsible, and the shape of the resultant spectra have been intensively studied in recent years with the aid of particle-in-cell (PIC) simulations (X. Li et al., 2021; F. Guo et al., 2024; L. Sironi et al., 2025). These simulations have enabled the study of kinetic reconnecting plasmas on increasingly large spatial and temporal scales.
@@
One of the key determinants of particle acceleration is the current sheet geometry: each particle acceleration mechanism requires certain configurations of electric and magnetic fields to operate. There are two broad categories of reconnecting current sheet geometries: single X-point and multiple X-point111also referred to as X-lines, the three-dimensional analogues of X-points. (H. Ji & W. Daughton, 2011). In the former, there is just one diffusion region where the field lines break and reconnect, while most of the reconnecting plasma proceeds directly from the inflow to the outflow exhaust. In the latter, the current sheet fragments due to tearing or plasmoid instability, resulting in a dynamic chain of magnetic islands with X-points in between (N. F. Loureiro et al., 2007). Note that in this paper we shall use X-point and diffusion region interchangeably to refer to areas where non-ideal electric fields dissipate inflowing oppositely-aligned magnetic flux.
Whether reconnection proceeds with a single X-point or with multiple X-points is critical to understanding particle acceleration dynamics. Recent work has established the importance of direct electric field acceleration at X-points in relativistic plasmas as an injection mechanism for subsequent acceleration (L. Sironi, 2022; S. R. Totorica et al., 2023; S. Gupta et al., 2025), a pathway which is more accessible in systems with more X-points. Furthermore, plasmoid-based Fermi acceleration mechanisms at later stages of energization (J. F. Drake et al., 2006; F. Guo et al., 2019; H. Hakobyan et al., 2021) are also primarily operative in the multiple X-point regime. In contrast, pickup acceleration occurs mainly in the single X-point regime for particles that do not pass through the X-point (J. F. Drake et al., 2009). How these mechanisms interact to produce the observed total spectra is a key research question.
@@
Given the strong dependence of particle acceleration on the current sheet geometry, it would be useful to have a framework to predict the transition from single X-point reconnection to multiple X-point reconnection. In the MHD model, this is accomplished by the Lundquist number $S\equiv Lv_{A}/(\eta c/4\pi)$, where $L$ is the half-length of the current sheet, $v_{A}$ is the Alfvén speed, and $\eta$ is the resistivity, assumed to be constant. When $S>S_{c}\sim 10^{4}$, the reconnecting current sheet fragments, resulting in fast plasmoid-mediated multiple X-point reconnection (H. Ji & W. Daughton, 2011). Of course, this fluid model does not always carry over to collisionless plasmas. Although recent strides have been made in understanding the tearing properties of collisionless current sheets (M. Hoshino, 2020; K. M. Schoeffler et al., 2024; K. Akutagawa et al., 2025), there is no comparable metric to a priori predict the global reconnection geometry in such collisionless systems, suggesting the need for kinetic studies.
@@
These questions also tie in closely to debates concerning the origin of fast reconnection. H. E. Petschek (1964) proposed that the slow Sweet-Parker reconnection (E. N. Parker, 1957) could be made fast by flanking a small diffusion region with four slow-mode shocks responsible for the majority of energy dissipation. His model has proven controversial since, in the absence of additional physics to localize the diffusion region, the Petschek solution is unstable and devolves to that of Sweet and Parker in MHD simulations (D. A. Uzdensky & R. M. Kulsrud, 2000). On the other hand, slow-mode shocks have been observed in the Earth’s magnetotail for decades (e.g., W. C. Feldman et al., 1984; Y. Saito et al., 1995; N. K. Walia et al., 2024), providing observational evidence in favor of Petschek’s theory. Large-domain kinetic simulations would thus be well suited to address whether Petschek-type outflows can form, and, if so, under what conditions.
@@
The use of electron-positron plasmas for this purpose is advantageous as it both enables large-scale simulations and facilitates the comparison of our results with the existing literature on high-$\sigma$ pair reconnection. However, we do expect our results to carry over to electron-ion plasmas, as we shall discuss later.
In this investigation, we focus on two magnetization values: $\sigma=0.3$ and $\sigma=10$, which have (upstream) Alfvén speeds $v_{A}=c\sqrt{\sigma/(\sigma+1)}$ of $\approx 0.48c$ and $\approx 0.95c$, respectively. The $\sigma=10$ case is well within the relativistic reconnection regime and, as we shall see, the resulting reconnection rate, current sheet structure, and particle energy spectra are consistent with the existing literature (e.g., L. Sironi et al., 2016; F. Guo et al., 2024). This acts as a control with which we may contrast the $\sigma=0.3$ case. Although we focus on only these two cases, we expect that the key results for $\sigma=10$ also apply to $\sigma>10$, and the results of $\sigma=0.3$ apply to $\sigma<0.3$, with the range $0.3<\sigma<10$ presenting the intermediate regime.
Our main finding is that the global structure of the current sheet is determined by the tug-of-war between the formation of wide-angle exhausts, which are naturally established to satisfy mass conservation, and plasmoid formation, which tends to perturb or disrupt those exhausts. If the number of plasmoids streaming forth from the diffusion regions is sufficiently large to prevent the exhausts from forming, the entire domain is thereby made available for multiple X-point reconnection and its associated mechanisms of particle acceleration (namely, direct $E$-field, Fermi, and betatron). However, if the plasmoid activity is insufficient to prevent the exhausts from being established, one is left with just a single X-point. Those canonical particle acceleration mechanisms are thereby throttled, yielding just a thermal Maxwellian distribution in particle energy.
This paper is organized as follows. In Section 2, we discuss the setup of our simulations. In Section 3, we discuss how the change in magnetization affects the global reconnection geometry. In Section 4, we discuss how those changes in the current sheet structure alter the particle acceleration pathways and resultant energy spectrum. In Section 5, we recapitulate our basic conclusions and discuss the broader implications of this work.
@@ 2 Simulation Setup
where $\Delta$ is set to 100 cells. Note that since no guide field is included, this is anti-parallel reconnection. The overdensity in the current sheet is initially $3n_{0}$. Reconnection is triggered by density depletion in the center of the current sheet, sending reconnection fronts outward in both directions. In the $x$-direction, fresh particles and flux are constantly supplied to replenish the reconnected plasma. These inflows are initialized cold, i.e., with $\theta\equiv T/mc^{2}=0$. In the $y$-direction, we use outflow boundary conditions which permit the reconnected plasma to be advected out of the system, as in H. Hakobyan et al. (2021). These outflow boundary conditions are crucial because they allow the system to reach steady-state equilibrium, a process which generally takes several Alfvén crossing times. This lengthy equilibration period is used to eliminate any sensitivity of our results to the details of the initial current sheet and to ensure that we are studying long-term behavior rather than initial transients.
We run simulations with a variety of spatial resolutions and domain sizes. The moderate-size and moderate-resolution Runs III and IV are used to illustrate the essential comparison between high- and low-$\sigma$. The small, highly resolved Runs I and II are used when considering effects at the scale of the diffusion region, such as X-point particle acceleration and the tearing rate within the diffusion region. Run V, covering the largest extent, is used to verify that the patterns observed in smaller subrelativistic simulations (Runs II and IV) carry over to very large system sizes. Ultimately, all simulations for each value of $\sigma$ show the same basic current sheet structure and particle spectrum – we do not observe any differences when varying the skin depth, implying that the higher resolutions used in Runs I-IV were unnecessary to capture the fundamental kinetic physics.
@@ 3.1 Petschek-like Exhausts
Reducing the magnetization from $\sigma=10$ to $\sigma=0.3$ results in a stark change in the structure of the current sheet, as can be seen in Figure 1. In the $\sigma=10$ case, as expected, the tearing instability results in copious plasmoid formation, frequent plasmoid mergers, and rapid plasmoid growth. The result is a self-similar current sheet made up of plasmoids of various sizes, as described in detail in L. Sironi et al. (2016). This is a clear case of multiple X-point reconnection wherein field lines reconnect at many locations along the current sheet. However, a primary X point can be identified at $y\approx 360\;c/\omega_{p}$ during this timestep.
@@
The situation for $\sigma=0.3$ is different. The current sheet near the X point at $y=240\;c/\omega_{p}$ is also tearing-unstable, resulting in the regular production of plasmoids. However, these plasmoids seldom merge and only a handful are present in the domain at any given time. The large-scale current sheet is a Petschek-like configuration: there is a small diffusion region spanning several skin depths in the center with exhausts at either end. By exhausts we refer to the moderate-density, weak-magnetic-field regions in the downstream, i.e., as distinguished from the plasmoids and the X-point(s).
The exhausts are even more pronounced in Run V that has a domain size ten times longer in the $y$-direction, shown in Figure 2. Here we see that when $L_{y}\gg c/\omega_{p}$, the plasmoids become nearly insignificant compared to the large laminar exhausts that form. While plasmoids do perturb these exhausts, they are too few in number to disrupt them: once the laminar exhausts stabilize, they persist for as long as we are able to simulate. Reconnection is still fast with the reconnection rate $R\equiv v_{\text{in}}/v_{A}\approx 0.10$, the same as for Run IV, and only somewhat lower than $R\approx 0.15$ for $\sigma=10$ (Run III).
Figure 2: Density profile for the $\sigma=0.3$ Run V spanning $6000\;c/\omega_{p}$ in the $y$-direction once it has reached steady state. Note that while plasmoids near the X-point have comparable widths to the width of the exhaust, by the time the plasmoids move farther away from the X-point their width becomes much smaller than that of the exhaust, and they become comparatively insignificant. Two plasmoids can be seen at $y\approx 1440$ and $1760\;c/\omega_{p}$. An animated version of this figure is available here.
The relative importance of plasmoids can be quantified by the plasmoid energy fraction, which we define as the quotient of the particle kinetic energy within plasmoids to the total particle kinetic energy within the overall current sheet. A region is considered part of a plasmoid if it is fully enclosed by a closed field line. The plot of the plasmoid energy fraction for each run is shown in Figure 3. Even in small runs that capture dynamics near the X-point, there is a clear difference between $\sigma=10$ runs where much more than 90% of the total energy is within plasmoids compared to less than 25% for $\sigma=0.3$. For $\sigma=0.3$, as the size of the domain increases, this fraction decreases while remaining roughly constant for the case $\sigma=10$.
Figure 3: Fraction of particle kinetic energy within plasmoids, considering just the current sheet. The rapid decline in plasmoid relevance for increasing system size for $\sigma=0.3$ reflects the growing dominance of the laminar exhausts. In the $\sigma=10$ case, with the exception of the diffusion regions, nearly all energy is stored within plasmoids. For several large simulations with $\sigma=10$ showing strong plasmoid dominance, see L. Sironi et al. (2016).
Thus, for low magnetization, the vast majority of energized particles avoid plasmoids altogether and simply proceeds directly from the cold inflow to the hot exhaust. At this interface slow-mode shocks form, as can be confirmed via comparison with the six Rankine-Hugoniot jump conditions representing continuity in mass, normal momentum, energy, normal magnetic field, transverse momentum, and tangential electric field, respectively (N. K. Walia et al., 2024):
@@
Here $\rho$ represents the mass density, $\mathbf{v}$ the flow velocity, $P_{g}$ the gas pressure, and $\gamma=5/3$ the adiabatic index. The components normal and tangential to the shock are denoted by $n$ and $t$, respectively. Figure 4 shows these quantities along a horizontal cut across the exhaust boundary, far from the X-point. As one can see, these jump conditions are well satisfied in our simulations, in general agreement with the Petschek model. Note, however, that, contrary to H. E. Petschek (1964), these slow-mode shocks are not switch-off shocks, consistent with earlier work (Y.-H. Liu et al., 2012). This is because current is spread throughout the exhaust rather than being localized only at the shock front, resulting in the field lines gradually curving in the exhaust rather than bending sharply (Figure 2). Across these slow-mode shocks, the magnetic pressure is converted into particle pressure and the density rises to $n/n_{0}\approx 3$. Mass conservation allows us to determine the approximate outflow angle $\theta$, as a function of this ratio and the reconnection rate $R\approx 0.1$:
@@
This results in an outflow angle of $\theta\approx 4^{\circ}$, consistent with the simulation (Figure 2).
Figure 4: Rankine-Hugoniot jump conditions for a cut along $y=1600\;c/\omega_{p}$ for Run V. The shock location is indicated by the green dashed line, the outflow is on the left, the inflow is on the right. Each conserved quantity is shown normalized relative to its upstream value.
Figure 5: Selected quantities across the same cut as in Figure 4, with the shock location indicated by the dotted line. The upstream magnetic pressure $P_{B0}\equiv B_{0}^{2}/8\pi$. As one would expect for a slow-mode shock, the density increases sharply whereas the magnetic field is greatly diminished. While the current peaks at the slow-mode shock, it does not vanish in the exhaust interior.

### v2

@@ 1 Introduction
Reconnection is a ubiquitous plasma process that rapidly converts magnetic energy into particle energy via rearrangement of the magnetic field topology. Originally posited to explain solar flares, reconnection now plays a key role in the theories of a multitude of astrophysical systems, from the heliosphere to compact objects. Observations and simulations suggest that reconnection can energize particles far beyond thermal energies, generating hard power-law energy spectra (L. Sironi & A. Spitkovsky, 2014; F. Guo et al., 2014). The conditions under which this particle acceleration occurs, the precise mechanisms responsible, and the shape of the resultant spectra have been intensively studied in recent years with the aid of particle-in-cell (PIC) simulations (X. Li et al., 2021; F. Guo et al., 2024; L. Sironi et al., 2025). These simulations have enabled the study of kinetic reconnecting plasmas on increasingly large spatial and temporal scales.
@@
One of the key determinants of particle acceleration is the current sheet geometry: each particle acceleration mechanism requires certain configurations of electric and magnetic fields to operate. There are two broad categories of reconnecting current sheet geometries: single X-point and multiple X-point111also referred to as X-lines, the three-dimensional analogues of X-points. (H. Ji & W. Daughton, 2011). In the former, there is just one diffusion region where the field lines break and reconnect, while most of the reconnecting plasma proceeds directly from the inflow to the outflow exhaust. In the latter, the current sheet fragments due to tearing or plasmoid instability, resulting in a dynamic chain of magnetic islands with X-points in between (N. F. Loureiro et al., 2007). [^2] In this paper we shall use X-point and diffusion region interchangeably to refer to areas where non-ideal electric fields dissipate inflowing oppositely-aligned magnetic flux. Recent work has established the importance of direct electric field acceleration at X-points in relativistic plasmas as an injection mechanism for subsequent acceleration (L. Sironi, 2022; S. R. Totorica et al., 2023; S. Gupta et al., 2025), a pathway which is more accessible in systems with more X-points. Furthermore, plasmoid-based Fermi acceleration mechanisms at later stages of energization (J. F. Drake et al., 2006; F. Guo et al., 2019; H. Hakobyan et al., 2021) are also primarily operative in the multiple X-point regime. In contrast, pickup acceleration occurs mainly in the single X-point regime for particles that do not pass through the X-point (J. F. Drake et al., 2009).
@@
Given the strong dependence of particle acceleration on the current sheet geometry, it would be useful to have a framework to predict the transition from single X-point reconnection to multiple X-point reconnection. In the MHD model, this is accomplished by the Lundquist number $S\equiv Lv_{A}/(\eta c/4\pi)$, where $L$ is the half-length of the current sheet, $v_{A}$ is the Alfvén speed, and $\eta$ is the resistivity, assumed to be constant. When $S>S_{c}\sim 10^{4}$, the reconnecting current sheet fragments, resulting in fast plasmoid-mediated multiple X-point reconnection (H. Ji & W. Daughton, 2011). Although recent strides have been made in understanding the tearing properties of collisionless current sheets (M. Hoshino, 2020; K. M. Schoeffler et al., 2024; K. Akutagawa et al., 2025), there is no comparable metric to a priori predict the global reconnection geometry in collisionless systems, suggesting the need for kinetic studies.
@@
These questions also tie in closely to debates concerning the origin of fast reconnection. H. E. Petschek (1964) proposed that the slow Sweet-Parker reconnection (E. N. Parker, 1957) could be made fast by flanking a small diffusion region with four slow-mode shocks responsible for the majority of energy dissipation. His model has proven controversial since, in the absence of additional physics to localize the diffusion region, the Petschek solution is unstable and devolves to that of Sweet and Parker in MHD simulations (D. A. Uzdensky & R. M. Kulsrud, 2000; T. G. Forbes et al., 2013). On the other hand, slow-mode shocks have been observed in the Earth’s magnetotail for decades (e.g., W. C. Feldman et al., 1984; Y. Saito et al., 1995; N. K. Walia et al., 2024), providing observational evidence in favor of Petschek’s theory. Large-domain kinetic simulations would thus be well suited to address whether Petschek-type outflows can form, and, if so, under what conditions.
@@
In this investigation, we focus on two magnetization values: $\sigma=0.3$ and $\sigma=10$, which have (upstream) Alfvén speeds $v_{A}=c\sqrt{\sigma/(\sigma+1)}$ of $\approx 0.48c$ and $\approx 0.95c$, respectively. The $\sigma=10$ case is well within the relativistic reconnection regime and, as we shall see, the resulting reconnection rate, current sheet structure, and particle energy spectra are consistent with the existing literature (e.g., L. Sironi et al., 2016; F. Guo et al., 2024). This acts as a control with which we may contrast the $\sigma=0.3$ case. Although we focus on only these two cases, we expect that the key results for $\sigma=10$ also apply to $\sigma>10$, and the results of $\sigma=0.3$ apply to $\sigma<0.3$, with the range $0.3<\sigma<10$ presenting an intermediate regime.
Our main finding is that the global structure of the current sheet is determined by the tug-of-war between the formation of wide-angle exhausts, which are naturally established to satisfy mass conservation, and plasmoid formation, which tends to perturb or disrupt those exhausts. If the number of plasmoids streaming forth from the diffusion regions is sufficiently large to prevent the exhausts from forming, the entire domain is thereby made available for multiple X-point reconnection and its associated mechanisms of particle acceleration (namely, direct $E$-field, Fermi, and betatron). However, if the plasmoid activity is insufficient to prevent the exhausts from being established, one is left with just a single X-point. The aforementioned canonical particle acceleration mechanisms are thereby throttled, yielding just a thermal Maxwellian distribution in particle energy.
This paper is organized as follows. In Section 2, we discuss the setup of our simulations. In Section 3, we discuss how the change in magnetization affects the global reconnection geometry. In Section 4, we discuss how those changes in the current sheet structure alter the particle acceleration pathways and resultant energy spectrum. In Section 5, we recapitulate our basic conclusions and discuss broader implications.
@@ 2 Simulation Setup
${\color[rgb]{0,0,0}{n=n_{0}\left(1+3\>\text{sech}^{2}{(x/\Delta)}\right)}},$ (3)
where $\Delta$ is set to 100 cells. Reconnection is triggered by density depletion in the center of the current sheet, sending reconnection fronts outward in both directions. In the $x$-direction, fresh particles and flux are constantly supplied to replenish the reconnected plasma. These inflows are initialized cold, i.e., with $\theta\equiv T/mc^{2}=0$. In the $y$-direction, we use outflow boundary conditions which permit the reconnected plasma to be advected out of the system (for details, see Appendix A). These outflow boundary conditions crucially allow the system to reach steady-state equilibrium, a process which generally takes several Alfvén crossing times. This lengthy equilibration period is used to eliminate any sensitivity of our results to the details of the initial current sheet and to ensure that we are studying long-term behavior rather than initial transients.
We run simulations with a variety of spatial resolutions and domain sizes, summarized in Table 1. The moderate-size and moderate-resolution Runs III and IV are used to illustrate the essential comparison between high- and low-$\sigma$. The small, highly resolved Runs I and II are used when considering effects at the scale of the diffusion region, such as X-point particle acceleration and the tearing rate within the diffusion region. Run V, covering the largest extent, is used to verify that the patterns observed in smaller subrelativistic simulations (Runs II and IV) carry over to very large system sizes. Ultimately, all simulations for each value of $\sigma$ show the same basic current sheet structure and particle spectrum—we do not observe any differences from varying the skin depth, implying that the higher resolutions used in Runs I-IV were unnecessary to capture the fundamental kinetic physics.
@@ 3.1 Petschek-like Exhausts
Reducing the magnetization from $\sigma=10$ to $\sigma=0.3$ results in a stark change in the structure of the current sheet, as can be seen in Figure 1. In the $\sigma=10$ case, as expected, the tearing instability results in copious plasmoid formation, frequent plasmoid mergers, and rapid plasmoid growth. The result is a self-similar current sheet made up of plasmoids of various sizes, as described in detail in L. Sironi et al. (2016). This is a clear case of multiple X-point reconnection wherein field lines reconnect at many locations along the current sheet. However, a primary X-point can be identified at $y\approx 360\;c/\omega_{p}$ during this timestep.
@@
The situation for $\sigma=0.3$ is different. The current sheet near the X-point at $y=240\;c/\omega_{p}$ is also tearing-unstable, resulting in the regular production of plasmoids. However, these plasmoids seldom merge and only a handful are present in the domain at any given time. The large-scale current sheet is a Petschek-like configuration: there is a small diffusion region spanning several skin depths in the center with exhausts at either end. By exhausts we refer to the moderate-density, weak-magnetic-field regions in the downstream, i.e., as distinguished from the plasmoids and the X-point(s).
The exhausts are even more pronounced in Run V that has a domain size ten times longer in the $y$-direction, shown in Figure 2. Here we see that when $L_{y}\gg c/\omega_{p}$, the plasmoids become nearly insignificant compared to the large laminar exhausts that form. While plasmoids do perturb these exhausts (as noted in M. E. Innocenti et al. (2015)), they are too few in number to disrupt them: once the laminar exhausts stabilize, they persist for as long as we are able to simulate. Reconnection is still fast with the reconnection rate $R\equiv v_{\text{in}}/v_{A}\approx 0.10$, the same as for Run IV, and only somewhat lower than $R\approx 0.15$ for $\sigma=10$ (Run III).
Figure 2: Density profile for the $\sigma=0.3$ Run V spanning $6000\;c/\omega_{p}$ in the $y$-direction once it has reached steady state. Note that while plasmoids near the X-point have comparable widths to the width of the exhaust, by the time the plasmoids move farther away from the X-point their width becomes much smaller than that of the exhaust, and they become comparatively insignificant. Two plasmoids can be seen at $y\approx 1440$ and $1760\;c/\omega_{p}$. An animated version of this figure shows the steady-state evolution over several Alfvén-crossing times.
Thus, for low magnetization, the vast majority of inflowing particles avoid plasmoids altogether and simply proceed directly from the cold inflow to the hot exhaust. At this interface slow-mode shocks form, as can be confirmed via comparison with the six Rankine-Hugoniot jump conditions representing continuity in mass, normal momentum, energy, normal magnetic field, transverse momentum, and tangential electric field, respectively (N. K. Walia et al., 2024):
@@
Here $\rho$ represents the mass density, $\mathbf{v}$ the flow velocity, $P_{g}$ the gas pressure, and $\gamma=5/3$ the adiabatic index. The components normal and tangential to the shock are denoted by $n$ and $t$, respectively. Figure 3 shows these quantities along a horizontal cut across the exhaust boundary, far from the X-point. As one can see, these jump conditions are well satisfied in our simulations, in general agreement with the Petschek model (see N. K. Walia et al. (2022) for a similar analysis of a hybrid simulation). Note, however, that, contrary to H. E. Petschek (1964), these slow-mode shocks are not also switch-off shocks, consistent with earlier work (Y.-H. Liu et al., 2012). This is because some current is spread throughout the exhaust rather than being localized only at the shock front (Fig 3), resulting in the field lines gradually curving in the exhaust rather than bending sharply (Figure 2). Across these slow-mode shocks, the magnetic pressure is converted into particle pressure and the density rises to $n/n_{0}\approx 3$. Mass conservation allows us to determine the approximate outflow angle $\theta$, as a function of this ratio and the reconnection rate $R\approx 0.1$:
@@
This results in an outflow angle of $\theta\approx 4^{\circ}$, consistent with the simulation (Figure 2). While the relativistic case (Fig. 1a) does not have this sort of persistent exhaust, one can note smaller “minijets” trailing the plasmoids as they advect out (J. Ortuño-Macías & K. Nalewajko, 2020). The minijets display a drop in tangential B field across their boundary similar to a slow-shock transition.
Figure 3: Plots (a) - (h) show selected quantities for a cut along $y=1600\;c/\omega_{p}$ in Run V. Plots (a) and (b) show the normalized density and normal bulk velocity, respectively. Plots (c) - (h) show the Rankine-Hugoniot jump quantities, normalized to their upstream values. Plot (i) shows general quantities plotted along that same $y$-value; the transition region depicted in plots (a) - (h) is shaded in gray. The upstream magnetic pressure $P_{B0}\equiv B_{0}^{2}/8\pi$.